In April, GGH Catering agreed to sell 1,800 ham & cheese sandwich trays to a private party. Each sandwich tray was priced at $44.00 and all of the trays were sold at that price. The contact person of the party told GGH management that they need more sandwich trays in May when the party is scheduled. Now, the private party needs 2,500 sandwich trays and the contact person has just agreed to buy each tray with a price of $30.00. By showing your calculations, decide of the price elasticity of this demand (elastic or inelastic)
Answer #1

Price elasticity of demand = % change in quantity demanded / % change in price

In this case, % change in quantity demanded = 2500-1800 / 1800 = 38.89%

% change in price = 30-44/44 = 31.82%

Thus, price elasticity of demand = 38.89/31.82 = 1.22

Since, it is >1 , it will be said that the demand for sandwiches is elastic.